# Stabilizing Rate of Stochastic Control Systems with Multiplicative Noise

### v1

@@ Abstract
This paper develops a quantitative framework for analyzing the mean-square exponential stabilization of stochastic linear systems with multiplicative noise, focusing specifically on the optimal stabilizing rate, which characterizes the fastest exponential stabilization achievable under admissible control policies. Our contributions are twofold. First, we extend norm-based techniques from deterministic switched systems to the stochastic setting, deriving a verifiable necessary and sufficient condition for the exact attainability of the optimal stabilizing rate, together with computable upper and lower bounds. Second, by restricting attention to state-feedback policies, we reformulate the optimal stabilizing rate problem as an optimal control problem with a nonlinear cost function and derive a Bellman-type equation. Since this Bellman-type equation is not directly tractable, we recast it as a nonlinear matrix eigenvalue problem whose valid solutions require strictly positive-definite matrices. To ensure the existence of such solutions, we introduce a regularization scheme and develop a Regularized Normalized Value Iteration (RNVI) algorithm, which in turn generates strictly positive-definite fixed points for a perturbed version of original nonlinear matrix eigenvalue problem while producing feedback controllers. Evaluating these regularized solutions further yields certified lower and upper bounds for the optimal stabilizing rate, resulting in a constructive and verifiable framework for determining the fastest achievable mean-square stabilization under multiplicative noise.
@@ 1 Introduction
In the theoretical framework of stochastic control, stabilizability analysis remains a core topic that permeates both fundamental research and engineering applications; classical theories focus centrally on a pivotal question for uncertain, noisy dynamic systems: can controller design guarantee bounded and asymptotically convergent trajectories? Classical results (e.g., H. J. Kushner (1967)) based on Lyapunov methods and Riccati theory provide affirmative answers under appropriate conditions. However, as modern engineering systems grow increasingly complex, merely proving stabilizability is no longer sufficient; the rate of stabilization, meaning how fast the state returns to the desired equilibrium, has become equally critical. This quantitative aspect directly influences both safety and performance in practical scenarios. For instance, in power and energy systems (Zhang and Cortés (2019); Weitenberg et al. (2018)), distributed frequency control has been shown to achieve exponential convergence and to examine how communication disruptions influence the resulting convergence rate, while transient frequency control enforces finite time return to a prescribed safe band with a guaranteed convergence rate, so the pace of recovery is treated as part of the design rather than a byproduct. For robotics and other safety-critical infrastructures (Ames et al. ; Nguyen and Sreenath (2016)), control Lyapunov and control barrier function methods often operate in tandem to enforce both stability and safety, and in particular, exponential barrier formulations embed explicit requirements on how rapidly safety constraints must be restored, indicating that the rate of convergence is an inherent component of reliable operation. Across these domains, stabilization feasibility alone is insufficient and one must also take into account how rapidly stability is achieved.
@@
Over the past decades, the classic theory of stability in deterministic systems is well established and provides the foundation for much of modern control. Lyapunov’s second method (Lyapunov (1992); Khalil and Grizzle (2002)) offers general certificates of asymptotic and exponential stability for nonlinear systems, while for linear time-invariant systems, algebraic Riccati equations and linear quadratic regulator (LQR) design yield constructive stabilizing state-feedback laws (Anderson and Moore (1971)). The development of linear matrix inequality (LMI) techniques (Boyd et al. (1994)) unifies and generalizes classic criteria, making stability analysis and controller synthesis numerically tractable. These methods constitute a mature toolbox: Lyapunov theory provides the conceptual basis, Riccati and LQR methods furnish constructive synthesis, and LMIs deliver computational scalability. Yet, all of these are primarily designed to decide whether stabilization is feasible and they rarely provide certified information about the rate of stabilization. When stochastic disturbances are present, the problem becomes significantly more challenging. In particular, systems with multiplicative noise, where randomness enters directly through system coefficients, arise naturally in applications ranging from portfolio optimization (Dombrovskii and Lyashenko (2003)) to systems biology (Yao et al. (2011)). For such systems, the standard notion is mean-square exponential stability (ESMS), which requires the second moment of the state to decay exponentially. This problem has been studied for decades, with early contributions dating back to the late 1960s (Kleinman (1969); McLane (1969); Wonham (1967)) and the continued development through numerous works Zhang et al. (2015); Feng et al. (2018a); Qi and Zhang (2018); Feng et al. (2018b); Li et al. (2021). Much of this literature establishes stability conditions via modified algebraic Riccati equations (Wonham (1967); Jacques and Jan (1976); Blankenship (1977)) or LMI formulations (Boyd et al. (1994); Elia (2005)), thereby connecting mean-square stability analysis to convex optimization. In parallel, a frequency-domain viewpoint is initiated by Willems and Blankenship (1971), who establishes necessary and sufficient conditions for mean-square stability in single-input single-output systems; this perspective has since been extended to multi-input multi-output cases (Hinrichsen and Pritchard (1996); Lu and Skelton (2002); Qi et al. (2017); Bamieh and Filo (2020)). From this standpoint, mean-square stability can be interpreted as a robust stability problem with respect to stochastic uncertainties. These approaches are powerful and constructive, but they remain qualitative: they tell us whether stabilization is possible, not how fast it can be achieved.
@@
To sum up, several important limitations remain in the existing literature. First, studies on stochastic systems with multiplicative noise provide tractable feasibility conditions, but they seldom address how fast such systems can be stabilized in the mean-square sense. Second, works on deterministic switched systems do quantify the fastest convergence rate, yet their methods rely critically on finite-dimensional properties absent in stochastic setting. Third, there remains no unified framework that can offer a computable and verifiable characterization of fastest stabilizing rate for stochastic systems with multiplicative disturbances. To bridge these gaps, this paper develops a unified framework for quantifying the optimal stabilizing rate, defined as the fastest mean-square exponential stabilization achievable under admissible control policies for stochastic systems with multiplicative noise. We first extend the norm-based approach of deterministic switched systems (Hu et al. (2017)) to the stochastic setting, thereby generalizing the quantitative notion of stabilization to random environments. Second, and more importantly, restricting to state-feedback policies, we reformulate the determination of fastest stabilizing rate as an optimal control problem with a nonlinear cost function. Building on this formulation, we propose the Regularized Normalized Value Iteration (RNVI) algorithm to compute certified upper and lower bounds for the fastest stabilizing rate under state-feedback control policies. The main contributions are summarized as follows.
@@
    This extension is technically nontrivial: norms on the infinite-dimensional space $L^{2}(\Omega;\mathbb{R}^{n})$ are no longer equivalent. We establish norm equivalence under suitable conditions and obtain a rigorous norm-based characterization in the stochastic setting.
@@
    To make this regularized approach constructive, we develop the RNVI algorithm. RNVI iteratively computes the positive-definite solutions and the corresponding feedback controllers for a sequence of decreasing regularization parameters. The procedure starts from a relatively large value of the regularization parameter, for which convergence is numerically reliable, and proceeds gradually through a continuation scheme that uses each computed solution as the initialization for the next stage. Each iteration provides verifiable upper and lower bounds for the optimal stabilizing rate under state-feedback control policies, and the tightest bounds obtain over all parameter values are reported as the final certified estimate.
Together, these results provide both theoretical guarantees and computational tools for analyzing and quantifying stochastic optimal stabilizing rate.
The remainder of the paper is organized as follows. Section 2 introduces the stochastic system model and the definition of the optimal stabilizing rate. Section 3 extends the norm-based framework of Hu et al. (2017) to our stochastic setting. Section 4 reformulates the problem as an optimal control problem and derives a Bellman-type equation which reduces to a nonlinear matrix eigenvalue equation. Section 5 develops the RNVI algorithm to obtain bounds for the optimal stabilizing rate. Section 6 reports numerical experiments that illustrate the performance of the proposed method. Section 7 concludes this paper.
@@ 2 Problem Formulation
Here, $x_{k}\in\mathbb{R}^{n}$ and $u_{k}\in\mathbb{R}^{m}$ are system state and control input, respectively. We assume that the initial value $x_{0}$ is deterministic. The noise sequence $\{\omega_{k}\}$ consists of independent and identically distributed (i.i.d) Gaussian random variables with zero mean and variance $\sigma^{2}$, and is assumed independent of $x_{0}$.
@@
is nonempty and is bounded below by 0; hence, its infimum $\rho(u)$ always exists. Since $\rho^{\ast}=\inf_{u\in\mathcal{U}}\rho(u)$ is the infimum of a nonempty set bounded below by 0, $\rho^{\ast}$ also exists. Our definition differs slightly from the deterministic setting of Hu et al. (2017). In Hu et al. (2017), the analogue of the growth inequality $\|x_{k}\|\leq\kappa\rho^{k}\|x_{0}\|$ used to define the stabilizing rate is written with an arbitrary vector norm on $\mathbb{R}^{n}$; in finite-dimensional deterministic systems all norms are equivalent, and the stabilizing rate is therefore independent of the chosen norm. For the stochastic system (1), the state evolves in $L^{2}(\Omega;\mathbb{R}^{n})$, where different norms are not equivalent in general. Consequently, the value of $\rho^{\ast}$ depends on the chosen norm. This paper works with the mean-square norm $\big(\mathbb{E}[x^{\top}x]\big)^{1/2}$.
@@ 3 Norm-Based Characterization of the Optimal Stabilizing Rate
We first establish the positive definiteness. For any $x_{0}$ and $u\in\mathcal{U}$, $\|x_{k}\|_{\mathrm{ms}}\geq 0$ implies $F(x_{0},u)\geq 0$ and hence $\eta(x_{0})\geq 0$. If $x_{0}=0$, choosing $u\equiv 0$ yields $x_{k}\equiv 0$ for all $k$, so $\eta(0)=0$. Conversely, if $x_{0}\neq 0$, then the term with $k=0$ gives $F(x_{0},u)\geq\|x_{0}\|_{\mathrm{ms}}>0$ for all $u$, and thus $\eta(x_{0})>0$.
@@
    There exist an admissible control policy $u\in\mathcal{U}$ and a constant $\kappa\geq 0$ such that $\mathbb{E}[x^{\top}_{k}x_{k}]\leq\kappa(\rho^{*})^{2k}\mathbb{E}[x_{0}^{\top}x_{0}]$ for all $k\in\mathbb{N}$.
@@
Hence, we know $\eta(x)<\infty$ for all $x$, $\mathcal{W}=L^{2}(\Omega;\mathbb{R}^{n})$, and $\sup_{\|x\|_{\mathrm{ms}}=1}\eta(x)\leq\sqrt{\kappa}<\infty.$ Thus (i) holds. ∎
While the necessary and sufficient condition ensuring exact attainability of $\rho^{\ast}$ is theoretically sound, it is also rather restrictive and may be difficult to verify in practice. Nevertheless, one can still derive meaningful and verifiable sufficient conditions that yield computable upper and lower bounds for $\rho^{\ast}$. Following the norm-based methodology of Hu et al. (2017), we now extend these bounding arguments to the stochastic setting considered here.
@@ 4 Optimal Control Reformulation of Problem (OSR)
Conversely, let $J(x_{0},u)=\alpha$, then for any $\varepsilon_{2}>0$ there exists $N$ such that for all $k>N$, it holds
@@
Since (7) and (8) hold in opposite directions for each admissible control $u\in\mathcal{U}$, we obtain the equality
@@
Taking the infimum over $u$ on both sides and using the monotonicity of $\log$ on $(0,\infty)$, we establish
@@
which is a subclass of the original admissible set $\mathcal{U}$. We emphasize that this restriction is imposed for technical reasons: working with feedback policies allows us to express all state trajectories in terms of a state-dependent performance function, which will be needed for the subsequent analysis. It is not claimed that feedback policies exhaust all admissible stabilizing policies; nevertheless, Ni et al. (2015) shows that for linear stochastic systems, open-loop $\ell^{2}$-stabilizable is equivalent to closed-loop $\ell^{2}$-stabilizable. Our use of feedback policies is therefore a structural restriction that facilitates the development of our framework, without making any claim about optimality of feedback relative to the general class $\mathcal{U}$. In particular, since $\mathcal{U}_{\mathrm{fb}}\subseteq\mathcal{U}$, the identity of Theorem 3 continues to hold when the admissible set is restricted to $\mathcal{U}_{\mathrm{fb}}$; namely,
@@
Suppose there exist a scalar $\lambda\in\mathbb{R}$ and a quadratic function $h(x)=x^{\top}Px$ with $P\in\mathbb{S}^{n}_{++}$ such that for every $x\in\mathbb{R}^{n}$

### v2

@@ Abstract
This paper develops a quantitative framework for analyzing the mean-square exponential stabilization of stochastic linear systems with multiplicative noise, focusing specifically on the optimal stabilizing rate, which characterizes the fastest exponential stabilization achievable under admissible control policies. Our contributions are twofold. First, we extend the norm-based techniques from deterministic switched systems to the stochastic setting, deriving a verifiable necessary and sufficient condition for the exact attainability of the optimal stabilizing rate, together with computable upper and lower bounds. Second, by restricting attention to state-feedback policies, we reformulate the optimal stabilizing rate problem as an optimal control problem with a nonlinear cost functional and derive a Bellman-type equation. Since this Bellman-type equation is not directly tractable, we recast it as a nonlinear matrix eigenvalue problem whose valid solutions require strictly positive-definite matrices. To ensure the existence of such solutions, we introduce a regularization scheme and develop a Regularized Normalized Value Iteration (RNVI) algorithm, which in turn converges to strictly positive-definite fixed points for a perturbed version of the original nonlinear matrix eigenvalue problem while producing feedback controllers. Evaluating these regularized solutions further yields certified lower and upper bounds for the optimal stabilizing rate, resulting in a constructive and verifiable framework for determining the fastest achievable mean-square stabilization under multiplicative noise.
@@ 1 Introduction
In the theoretical framework of stochastic control, stabilizability analysis remains a core topic that permeates both fundamental research and engineering applications; classic theories focus centrally on a pivotal question for uncertain, noisy dynamic systems: can controller design guarantee bounded and asymptotically convergent trajectories? Classic results (e.g., H. J. Kushner (1967)) based on Lyapunov methods and Riccati theory provide affirmative answers under appropriate conditions. However, as modern engineering systems grow increasingly complex, merely proving stabilizability is no longer sufficient; the rate of stabilization, meaning how fast the state returns to the desired equilibrium, has become equally critical. This quantitative aspect directly influences both safety and performance in practical scenarios. For instance, in power and energy systems (Zhang and Cortés (2019); Weitenberg et al. (2018)), distributed frequency control has been shown to achieve exponential convergence and to examine how communication disruptions influence the resulting convergence rate, while transient frequency control enforces finite time return to a prescribed safe band with a guaranteed convergence rate, and thus the pace of recovery is treated as part of the design rather than a byproduct. For robotics and other safety-critical infrastructures (Ames et al. ; Nguyen and Sreenath (2016)), control Lyapunov and control barrier function methods often operate in tandem to enforce both stability and safety, and in particular, exponential barrier formulations embed explicit requirements on how rapidly safety constraints must be restored, indicating that the rate of convergence is an inherent component of reliable operation. Across these domains, stabilization feasibility alone is insufficient and one must also take into account how rapidly stability is achieved.
@@
Over the past decades, the classic theory of stability in deterministic systems is well established and provides the foundation for much of modern control. Lyapunov’s second method (Khalil and Grizzle (2002)) offers general certificates of asymptotic and exponential stability for nonlinear systems, while for linear time-invariant systems, algebraic Riccati equations and linear quadratic regulator (LQR) design yield constructive stabilizing state-feedback laws (Anderson and Moore (1971)). The development of linear matrix inequality (LMI) technique (Boyd et al. (1994)) unifies and generalizes classic criteria, making stability analysis and controller synthesis numerically tractable. These methods constitute a mature toolbox: Lyapunov theory provides the conceptual basis, Riccati and LQR methods furnish constructive synthesis, and LMIs deliver computational scalability. Yet, all of these are primarily designed to decide whether stabilization is feasible and they rarely provide certified information about the rate of stabilization. When stochastic disturbances are present, the problem becomes significantly more challenging. In particular, systems with multiplicative noise, where randomness enters directly through system coefficients, arise naturally in applications ranging from portfolio optimization (Dombrovskii and Lyashenko (2003)) to systems biology (Yao et al. (2011)). For such systems, the standard notion is mean-square exponential stability (ESMS), which requires the second moment of the state to decay exponentially. This problem has been studied for decades, with early contributions dating back to the late 1960s (Kleinman (1969); McLane (1969); Wonham (1967)) and the continued development through numerous works Zhang et al. (2015); Feng et al. (2018a); Qi and Zhang (2018); Feng et al. (2018b); Li et al. (2021). Much of this literature establishes stability conditions via modified algebraic Riccati equations (Wonham (1967); Jacques and Jan (1976); Blankenship (1977)) or LMI formulations (Boyd et al. (1994); Elia (2005)), thereby connecting mean-square stability analysis to convex optimization. In parallel, a frequency-domain viewpoint is initiated by Willems and Blankenship (1971), who establish necessary and sufficient conditions for mean-square stability in single-input single-output systems; this perspective has since been extended to multi-input multi-output cases (Hinrichsen and Pritchard (1996); Lu and Skelton (2002); Qi et al. (2017); Bamieh and Filo (2020)). From this standpoint, mean-square stability can be interpreted as a robust stability problem with respect to stochastic uncertainties. These approaches are powerful and constructive, but they remain qualitative: they tell us whether stabilization is possible, not how fast it can be achieved.
@@
To sum up, several important limitations remain in the existing literature. First, studies on stochastic systems with multiplicative noise provide tractable feasibility conditions, but they seldom address how fast such systems can be stabilized in the mean-square sense. Second, works on deterministic switched systems do quantify the fastest convergence rate, yet their methods rely critically on finite-dimensional properties that are absent in the stochastic setting. Third, there remains no unified framework that can offer a computable and verifiable characterization of fastest stabilizing rate for stochastic systems with multiplicative disturbances. To bridge these gaps, this paper develops a unified framework for quantifying the optimal stabilizing rate, defined as the fastest mean-square exponential stabilization achievable under admissible control policies for stochastic systems with multiplicative noise. We first extend the norm-based approach of deterministic switched systems (Hu et al. (2017)) to the stochastic setting, thereby generalizing the quantitative notion of stabilization to random environments. Second, and more importantly, restricting to state-feedback policies, we reformulate the determination of fastest stabilizing rate as an optimal control problem with a nonlinear cost functional. Building on this formulation, we propose the Regularized Normalized Value Iteration (RNVI) algorithm to compute the certified upper and lower bounds for the fastest stabilizing rate under state-feedback control policies. The main contributions are summarized as follows.
@@
    This extension is technically nontrivial: norms on the infinite-dimensional space $L^{2}(\Omega;\mathbb{R}^{n})$ are no longer equivalent. We establish some norm equivalence under suitable conditions and obtain a rigorous norm-based characterization in the stochastic setting.
@@
    To make this regularized approach constructive, we develop the RNVI algorithm. RNVI iteratively computes the positive-definite solutions and the corresponding feedback controllers for a sequence of decreasing regularization parameters. The procedure starts from a relatively large value of the regularization parameter, for which the convergence is numerically reliable, and proceeds gradually through a continuation scheme that uses each computed solution as the initialization for the next stage. Each iteration provides verifiable upper and lower bounds for the optimal stabilizing rate under state-feedback control policies, and the tightest bounds obtain over all parameter values are reported as the final certified estimate.
Together, these results provide both theoretical guarantees and computational tools for analyzing and quantifying the stochastic optimal stabilizing rate.
The remainder of the paper is organized as follows. Section 2 introduces the stochastic system model and the definition of the optimal stabilizing rate. Section 3 extends the norm-based framework of Hu et al. (2017) to our stochastic setting. Section 4 reformulates the problem as an optimal control problem and derives a Bellman-type equation, which is shown to reduce to a nonlinear matrix eigenvalue equation. Section 5 develops the RNVI algorithm to obtain bounds for the optimal stabilizing rate. Section 6 reports numerical experiments that illustrate the performance of the proposed method. Section 7 then concludes this paper.
@@ 2 Problem Formulation
Here, $x_{k}\in\mathbb{R}^{n}$ and $u_{k}\in\mathbb{R}^{m}$ are the system state and control input, respectively. We assume that the initial state $x_{0}$ is Gaussian with mean $\bar{x}_{0}$ and covariance $\Sigma_{\mathrm{init}}$. The noise sequence $\{\omega_{k}\}$ consists of independent and identically distributed (i.i.d) Gaussian random variables with zero mean and variance $\sigma^{2}$, and is assumed independent of $x_{0}$.
@@
is nonempty and is bounded below by 0; hence, its infimum $\rho(u)$ always exists. Since $\rho^{\ast}=\inf_{u\in\mathcal{U}}\rho(u)$ is the infimum of a nonempty set bounded below by 0, $\rho^{\ast}$ also exists. Our definition differs slightly from the deterministic setting of Hu et al. (2017). In Hu et al. (2017), the analogue of the growth inequality $\|x_{k}\|\leq\kappa\rho^{k}\|x_{0}\|$ used to define the stabilizing rate is written with an arbitrary vector norm on $\mathbb{R}^{n}$; in finite-dimensional deterministic systems, all norms are equivalent and the stabilizing rate is therefore independent of the chosen norm. For the stochastic system (1), the state evolves in $L^{2}(\Omega;\mathbb{R}^{n})$, where different norms are not equivalent in general. Consequently, the value of $\rho^{\ast}$ depends on the chosen norm. This paper works with the mean-square norm $\big(\mathbb{E}[x^{\top}x]\big)^{1/2}$.
@@ 3 Norm-Based Characterization of the Optimal Stabilizing Rate
We first establish the positive definiteness. For any $x_{0}$ and $u\in\mathcal{U}$, $\|x_{k}\|_{\mathrm{ms}}\geq 0$ implies $F(x_{0},u)\geq 0$ and hence $\eta(x_{0})\geq 0$. If $x_{0}=0$, choosing $u\equiv 0$ yields $x_{k}\equiv 0$ for all $k$, and thus $\eta(0)=0$. Conversely, if $x_{0}\neq 0$, then the term with $k=0$ gives $F(x_{0},u)\geq\|x_{0}\|_{\mathrm{ms}}>0$ for all $u$, and thus $\eta(x_{0})>0$.
@@
    There exist an admissible control policy $u\in\mathcal{U}$ and a constant $\kappa\geq 0$ such that $\mathbb{E}[x^{\top}_{k}x_{k}]\leq\kappa(\rho^{*})^{2k}\mathbb{E}[x_{0}^{\top}x_{0}]$ holds for all $k\in\mathbb{N}$.
@@
Hence, we know $\eta(x)<\infty$ for all $x$, $\mathcal{W}=L^{2}(\Omega;\mathbb{R}^{n})$, and $\sup_{\|x\|_{\mathrm{ms}}=1}\eta(x)\leq\sqrt{\kappa}<\infty.$ Thus, (i) holds. ∎
While the necessary and sufficient condition ensuring exact attainability of $\rho^{\ast}$ is theoretically sound, it is also rather restrictive and may be difficult to be verified in practice. Nevertheless, one can still derive meaningful and verifiable sufficient conditions that yield computable upper and lower bounds for $\rho^{\ast}$. Following the norm-based methodology of Hu et al. (2017), we now extend these bounding arguments to the stochastic setting considered here.
@@ 4 Optimal Control Reformulation of Problem (OSR)
Conversely, let $J(x_{0},u)=\alpha$, then for any $\varepsilon_{2}>0$ there exists $N$ such that for all $k>N$ it holds
@@
Since (7) and (8) hold in the opposite directions for each admissible control $u\in\mathcal{U}$, we obtain the equality
@@
Taking the infimum over $u$ on both sides and using the monotonicity of function $\log$ on $(0,\infty)$, we establish
@@
where $\{x_{k}\}$ denotes the state trajectory of system (1) driven by $u$, and $\mathcal{U}_{\mathrm{fb}}$ is a subclass of the original admissible set $\mathcal{U}$. We emphasize that this restriction is imposed for technical reasons: working with feedback policies allows us to express all state trajectories in terms of a state-dependent performance function, which will be needed for the subsequent analysis. It is not claimed that feedback policies exhaust all admissible stabilizing policies; nevertheless, Ni et al. (2015) shows that for linear stochastic systems, open-loop $\ell^{2}$-stabilizable is equivalent to closed-loop $\ell^{2}$-stabilizable. Our use of feedback policies is therefore a structural restriction that facilitates the development of our framework, without making any claim about optimality of feedback relative to the general class $\mathcal{U}$. In particular, since $\mathcal{U}_{\mathrm{fb}}\subseteq\mathcal{U}$, the identity of Theorem 3 continues to hold when the admissible set is restricted to $\mathcal{U}_{\mathrm{fb}}$; namely,
@@
Suppose there exist a scalar $\lambda\in\mathbb{R}$ and a quadratic function $h(x)=x^{\top}Px$ with $P\in\mathbb{S}^{n}_{++}$ such that for every $x\in\mathbb{R}^{n}$,